OpenStudy (anonymous):

graph -x - 2y = 6. can anybody do this?!

OpenStudy (anonymous):

Try to do that for youself. If you cant then ask

OpenStudy (anonymous):

thats a line that intersects y axis at -3 and has a negative inclination of 0.5

OpenStudy (anonymous):

i dont understand how to do it.

OpenStudy (anonymous):

im trying not to give you the answer but help you understanding the method instead. Think with me, lets solve it for Y -2y = x +6, right? Now lets get rid of the "minus" signal. Multiply both members of equation by (-1) 2y = -x -6 Not devide both members by (2) y = -0.5 -3 now you have it in the y = M.x + B format. Where M is the inclination = -0.5 and B = y axis intersection = -3 Now draw the point (0, -3) And now you can give a number to X and find its Y, example: If X = 2, then: y = 0.5 * 2 -3 y = -2 Now draw the point (2, -2) if X = 4, then: y = 0.5 * 4 -3 y = 2-3 = -1 Now draw the point (4, -1) and now that you have 3 points, just draw a line between them